# Look Twice before You Leap: A Rational Agent Framework for Localized Adversarial Anonymization

### v1

@@ Abstract
Current LLM-based text anonymization frameworks usually rely on remote API services from powerful LLMs, which creates an inherent "privacy paradox": users must somehow disclose data to untrusted third parties for superior privacy preservation. Moreover, directly migrating these frameworks to local small-scale models (LSMs) offers a suboptimal solution with catastrophic collapse in utility based on our core findings. Our work argues that this failure stems not merely from the capability deficits of LSMs, but from the inherent irrationality of the greedy adversarial strategies employed by current state-of-the-art (SoTA) methods. We model the anonymization process as a trade-off between Marginal Privacy Gain (MPG) and Marginal Utility Cost (MUC), and demonstrate that greedy strategies inevitably drift into an irrational state. To address this, we propose Rational Localized Adversarial Anonymization (RLAA), a fully localized and training-free framework featuring an Attacker-Arbitrator-Anonymizer (A-A-A) architecture. RLAA introduces an arbitrator that acts as a rationality gatekeeper, validating the attacker’s inference to filter out feedback providing negligible benefits on privacy preservation. This mechanism enforces a rational early-stopping criterion, and systematically prevents utility collapse. Extensive experiments on different datasets demonstrate that RLAA achieves the best privacy utility trade-off, and in some cases even outperforms SoTA on the Pareto principle. Our code and datasets will be released upon acceptance.
@@
Donghang Duan    Xu Zheng University of Electronic Science and Technology of China
@@ 1 Introduction
With the proliferation of large language models (LLMs) adopted in various fields, processing textual data that potentially contain sensitive Personal Identifiable Information (PII) has become inevitable and ubiquitous deusser2025survey. This trend makes text anonymization crucial, not only to meet stringent regulatory requirements such as GDPR and CCPA albrecht2016gdpr; bonta2022california, but also to protect personal privacy while unlocking the value of data. However, this situation presents a tricky trade-off between semantic utility and privacy where over-anonymization destroys the semantic value while under-anonymization invites re-identification risks, which is particularly acute in sensitive domains like smart healthcare and legal technology morris2025diri; im2024exploring; liu2025legal. The rapid evolution of LLM has exacerbated this situation because of their dual roles as context-aware anonymizers and superior attackers staab2024large; wang2025unique. Consequently, traditional NER-based anonymization methods are rendered obsolete, shifting the paradigm towards complex semantic obfuscation that conceals latent identity with less sacrifice of utility loiseau2025tau.
@@
To address this situation, advanced anonymization frameworks based on LLMs have emerged. As the current state-of-the-art (SoTA) method, Feedback-guided Adversarial Anonymization (FgAA) staab2024large employs a dynamic adversarial game: an attacker model attempts to re-identify personal information from the text generated by an anonymizer model, and the anonymizer uses the inference feedback to instruct the anonymization process for refinement. However, the superior performance of FgAA heavily relies on the capabilities of powerful or closed-source LLMs (e.g., GPT-4 achiam2023gpt or DeepSeek-V3.2-Exp deepseekai2024deepseekv32 in our research), which are typically accessible only via third-party APIs. This inherent dependency creates a fundamental "privacy paradox": to protect sensitive data, users must first send them to an external and uncontrolled commercial entity, which constitutes a critical privacy vulnerability smith2023identifying. This problem causes trust failure feretzakis2024trustworthy, and renders the architecture unacceptable in high-sensitivity scenarios.
Figure 1: Utility Failure of FgAA’s Naive Migration.
@@
The most intuitive solution to this privacy paradox is to migrate adversarial frameworks to local small-scale models (LSMs). However, our empirical findings reveal that this naive migration results in a catastrophic utility collapse. As illustrated in Figure 1, the naive migration on Llama3-8B causes destructive over-editing. It indiscriminately strips away non-sensitive context and stylistic nuances, reducing the rich original narrative to a generic and vacuous summary ("The days are still memorable"). Besides, it is observed that this collapse persists even when implementing instruction refinement (e.g., outputting "unknown" for inferences with low confidence), which demonstrates that heuristic constraints alone cannot rectify the inherent irrationality of the greedy decision-making process.
@@
We argue that this failure stems not merely from the capability deficits of LSMs, but from the inherent irrationality of the greedy adversarial strategies employed by current methods. Adopting an economic perspective, the anonymization process constitutes a trade-off between Marginal Privacy Gain (MPG) and Marginal Utility Cost (MUC). By quantifying this ratio as the Marginal Rate of Substitution ($\text{MRS}=\frac{\text{MUC}}{\text{MPG}}$), the root cause of the utility collapse becomes mathematically evident: Greedy strategies inevitably drift into an irrational state due to LSMs’ hallucinations and the law of diminishing returns, where the model executes destructive edits for non-existent or negligible leaks (see theoretical proofs in Appendix A). In this case, MUC continues to accumulate while MPG approaches zero, pushing the MRS to inefficient levels. A concurrent study called SEAL kim2025self aims to distill knowledge from a powerful LLM to an LSM by using Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) rafailov2023direct to avoid disclosing raw data to third-party APIs. However, it requires additional resources for post-training and high-quality synthetic datasets remain scarce and difficult to construct effectively yukhymenko2024synthetic. Besides, our experimental results in Section 4.2 shows that it still fails to handle the rationality of anonymization.
@@
To address the irrationality of naive migration and the data scarcity of distillation methods, we propose RLAA (Rational Localized Adversarial Anonymization), a training-free framework designed for fully localized deployment. The core idea of RLAA is to compensate for the capability deficits of LSMs through architectural rationality rather than finetuning. RLAA features a novel "Attacker-Arbitrator-Anonymizer" (A-A-A) architecture, where the arbitrator functions as a rationality gatekeeper. Instead of blindly accepting feedback, this component compels the system to scrutinize the attacker’s suggestions before executing text modifications. Drawing on the idea of meta-reasoning alazraki2025meta; wen2025reinforcement, the arbitrator implements a dual-validation mechanism: it first filters hallucinations to ensure factual correctness, and simultaneously enforces a rational stopping criterion based on the marginal rate of substitution (MRS). Crucially, this mechanism prevents utility collapse by rejecting destructive edits driven by invalid feedback. Furthermore, it maintains robust privacy protection by focusing the modification budget solely on genuine leaks, effectively eliminating the deadweight loss of irrational editing without compromising valid defense. We conducted a comprehensive evaluation spanning baseline comparisons, component ablations, and generalization stress tests across mainstream LLMs (Llama3-8B llama3modelcard, Qwen2.5-7B qwen2.5) and DeepSeek-V3.2-Exp deepseekai2024deepseekv32. Additionally, we validated the agent’s rational decision-making process through a quantitative economic efficiency analysis based on the MRS metric. The results empirically demonstrate that RLAA structurally prevents utility collapse and achieves a superior privacy-utility trade-off compared to our baselines.
In summary, the main contributions of this study are as follows:
@@
  • It is argued that migrating adversarial anonymization to local environments is crucial for eliminating API-based privacy risks, but our research shows that a naive migration to LSMs results in utility failure. This failure is analyzed and attributed to the inherent irrationality of greedy adversarial strategies via a marginal economic analysis in our work.
@@
  • RLAA is proposed as a localized and training-free framework featuring an Attacker-Arbitrator-Anonymizer (A-A-A) architecture. By introducing an arbitrator inspired by meta-reasoning, it structurally enforces rational decision-making to prevent utility collapse.
@@
  • Extensive experiments demonstrate that RLAA achieves a superior privacy-utility trade-off compared to our baselines. Notably, it achieves strict Pareto dominance on the reddit-self-disclosure dataset and demonstrate its mechanism generalization across different base models.
@@ 2.1 Inference Attacks through LLMs
With increased reasoning capabilities, LLMs have evolved into potent privacy attackers capable of automatically inferring personal attributes from unstructured text staab24beyond; wang2025unique. staab24beyond demonstrated that models can accurately deduce fine-grained PII from casual online comments. Extending to high-stakes domains, nyffenegger2024anonymity revealed that LLMs could re-identify individuals in court decisions. Recent audits by panda2025daiq further confirmed that this inference capability is consistent across various model architectures. These studies show that LLM significantly lowers the threshold for privacy violations, enabling re-identification attacks to occur on an unprecedented scale staab24beyond.
@@ 2.2 Text Anonymization
In recent years, the studies on text anonymization has evolved from statistical obfuscation to semantic rewriting, and we categorize related works into the following categories.
Traditional Approaches.
In early efforts, Differential Privacy (DP) methods abadi2016deep; wu2023privacy; igamberdiev2023dp and representation learning coavoux2018privacy; friedrich2019adversarial served as the standard. However, these methods often compromise textual readability or require training specific encoders, which limits their applicability to open-ended generation tasks. To address utility preservation, early adversarial frameworks emerged: shetty2018a4nt introduced A4NT to obfuscate author attributes via adversarial NMT, while mosallanezhad2019deep utilized Deep Reinforcement Learning to optimize the privacy-utility trade-off. These works laid the foundational game mechanic for modern anonymizers.
LLM-driven Adversarial Frameworks.
With the rise of generative AI, Feedback-guided Adversarial Anonymization (FgAA) staab2024large has become the representative state-of-the-art. It systematizes the LLM’s duality, leveraging an attacker model to iteratively critique and refine the anonymizer’s output. Following this paradigm, SEAL kim2025self attempts to distill these capabilities into small models via SFT and DPO. However, this distillation mechanism remains constrained by the scarcity of high-quality synthetic datasets and potential generalization issues. Beyond the standard rewriting loop of FgAA, other frameworks explore distinct mechanisms. IncogniText frikha2024incognitext employs a "misleading" strategy: instead of simply removing PII, it injects false attributes to randomize the attacker’s inference. RUPTA yang2024robust focuses on downstream utility (e.g., occupation classification), employing specialized evaluators to balance privacy against specific task performance. Crucially, a common limitation across these advanced frameworks is their reliance on strong LLMs via API or extensive training pipelines, necessitating the migration to fully local and general solutions.
Figure 2: Overview of the RLAA Framework. The arbitrator serves as a rationality gatekeeper within the adversarial loop. It validates and filters the attacker’s feedback to ensure that the anonymizer only performs cost-effective modifications.
@@ 3.1 Threat Model
The proposed framework is designed to defend against two distinct adversaries in the text anonymization task.
Service Provider Adversary ($\mathcal{A}_{service}$).
This adversary represents the third-party entity hosting remote LLM services. It embodies the "privacy paradox": to utilize superior anonymizers, users must first transmit raw sensitive text $x^{ori}$ to the provider. This transmission exposes users to risks such as unauthorized data retention for model training or commercial profiling smith2023identifying. Distinct from probable inference attacks, we characterize this threat as a deterministic data exposure, where $\mathcal{A}_{service}$ gains full access to $x^{ori}$ immediately upon submission. Our framework can structurally eliminate this threat through fully local deployment.
Re-identification Adversary ($\mathcal{A}_{re-id}$).
This adversary represents any entity with access to the anonymized output $x^{ano}$. We assume $\mathcal{A}_{re-id}$ employs a powerful LLM $\mathcal{M}_{attack}$ (e.g., DeepSeek-V3.2-Exp in our study) to infer Personal Identifiable Information (PII). Formally, given an attack prompt $P$, the adversary’s goal is to maximize the accuracy of the inferred attribute-value tuples $\{(a_{j},v_{j})\}$, where $a_{j}$ denotes a specific private attribute category (e.g., location) and $v_{j}$ represents its corresponding inferred value:
$\{(a_{j},v_{j})\}\leftarrow\mathcal{M}_{attack}(P(x^{ano}))$ (1)
@@ 3.2 Problem Formulation
The text anonymization task can be formalized as a constrained transformation problem. Let $\mathcal{X}$ be the space of all text sequences. Given an original text $x^{(0)}\in\mathcal{X}$, let $\mathcal{A}_{priv}=\{(a_{k},v_{k}^{*})\}_{k=1}^{K}$ be the set of ground-truth private attribute-value pairs contained in $x^{(0)}$, where $a_{k}$ represents the attribute category (e.g., age) and $v_{k}^{*}$ denotes its true value.
@@
$P_{atk}(x)=\frac{1}{K}\sum_{k=1}^{K}\mathbb{I}\left(\hat{v}_{k}=v_{k}^{*}\mid(a_{k},\hat{v}_{k})\in\mathcal{M}_{atk}(x)\right)$ (2)
where $\mathbb{I}(\cdot)$ is the indicator function checking if the inferred value $\hat{v}_{k}$ matches the ground truth $v_{k}^{*}$. Let $U(x^{(0)},x):\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be a utility function measuring semantic preservation, where a higher value indicates better utility.
The ideal objective is to find an optimum $x^{*}$ such that private attributes are protected while utility is maximized:
$\max_{x}U(x^{(0)},x)\quad\text{s.t.}\quad P_{atk}(x)\leq\delta$ (3)
where $\delta$ represents the maximum acceptable privacy risk.
Migrating the framework to a local environment naturally eliminates the service provider threat $\mathcal{A}_{service}$. However, this introduces a critical challenge: directly optimizing Eq. 3 via greedy adversarial iterations with LSMs often leads to utility collapse. We hypothesize that this failure arises because naive greedy agents lack a mechanism to evaluate the cost-effectiveness of privacy gains, leading to irrational modifications where the utility cost outweighs the privacy benefit.
To rigorously address this, we reframe the iterative process over steps $t=1,\dots,T$ through an economic lens, viewing each anonymization operation as a transaction. We define the following marginal metrics to quantify the rationality of each update:
@@
$\Delta P_{t}=P(A_{priv}|x^{(t-1)})-P(A_{priv}|x^{(t)})$ (4)
@@
The instantaneous "price" of privacy preservation, representing the utility cost per unit of privacy gained:
@@
From this perspective, a rational agent must operate under the constraint $\text{MRS}_{t}\leq\lambda$, where $\lambda$ represents the maximum acceptable utility cost per unit of privacy. Naive adversarial agents typically violate this condition, drifting into irrational states of deadweight loss (detailed in Appendix A). Guided by this insight, a rationality-constrained framework is proposed to enforce efficient optimization.

### v2

@@ Abstract
Current LLM-based text anonymization frameworks usually rely on remote API services from powerful LLMs, which creates an inherent privacy paradox: users must disclose data to untrusted third parties for guaranteed privacy preservation. Moreover, directly migrating current solutions to local small-scale models (LSMs) offers a suboptimal solution with severe utility collapse. Our work argues that this failure stems not merely from the capability deficits of LSMs, but significantly from the inherent irrationality of the greedy adversarial strategies employed by current state-of-the-art (SOTA) methods. To address this, we propose Rational Localized Adversarial Anonymization (RLAA), a fully localized and training-free framework featuring an Attacker-Arbitrator-Anonymizer architecture. We model the anonymization process as a trade-off between Marginal Privacy Gain (MPG) and Marginal Utility Cost (MUC), and demonstrate that greedy strategies tend to drift into an irrational state. Instead, RLAA introduces an arbitrator that acts as a rationality gatekeeper, validating the attacker’s inference to filter out feedback providing negligible privacy benefits. This mechanism promotes a rational early-stopping criterion, and structurally prevents utility collapse. Extensive experiments on different benchmarks demonstrate that RLAA achieves a superior privacy-utility trade-off compared to strong baselines.
@@
Donghang Duan1, Xu Zheng1, Yuefeng He1, Chong Mu1, Leyi Cai2, Lizong Zhang1 1School of Computer Science and Engineering, University of Electronic Science and Technology of China 2School of Information, Renmin University of China
@@ 1 Introduction
Large language models (LLMs) are increasingly deployed to process real-world text containing sensitive Personal Identifiable Information (PII), such as medical records, legal documents and online self-disclosures bommasani2021opportunities; weidinger2021ethical; deusser2025survey. To comply with stringent regulatory requirements like GDPR and CCPA albrecht2016gdpr; bonta2022california, effective text anonymization has become a prerequisite for the responsible use of such data. However, this situation presents a tricky trade-off between semantic utility and privacy, where over-anonymization destroys the semantic value while under-anonymization invites re-identification risks, which is particularly acute in sensitive domains like smart healthcare and legal technology im2024exploring; morris2025diri; liu2025legal. The rapid evolution of LLMs has exacerbated this situation because of their dual roles as context-aware anonymizers and superior attackers wang2025unique. As a result, traditional NER-based anonymization methods are increasingly inadequate dernoncourt2017identification, motivating a shift toward semantic anonymization that conceals latent identity cues beyond surface-level entities loiseau2025tau.
@@
To address this situation, advanced anonymization frameworks based on LLMs have emerged. As the current state-of-the-art (SOTA) framework, Feedback-guided Adversarial Anonymization (FgAA) staab2024large typically employs a dynamic adversarial game: an attacker model attempts to re-identify personal information from the text generated by an anonymizer model, and the anonymizer uses the inference feedback to instruct the anonymization process for refinement. Despite its strong empirical performance, FgAA heavily relies on the capabilities of powerful or closed-source LLMs like GPT-4 achiam2023gpt and DeepSeek-V3.2-Exp deepseekai2024deepseekv32, which are typically accessible only via third-party APIs. This architectural dependency gives rise to a fundamental privacy paradox: to anonymize sensitive data, users must first transmit the raw content to external and uncontrollable service providers, which constitutes a critical privacy vulnerability smith2023identifying and renders the architecture unacceptable in high-sensitivity scenarios feretzakis2024trustworthy. Although subsequent works like IncogniText frikha2024incognitext and SEAL kim2025self aim to mitigate this risk through localized deployment, they compromise semantic utility by injecting ungrounded hallucinations or require complex training pipelines.
Figure 1: Utility Collapse of FgAA’s Naive Migration.
@@
A natural and simple solution to the privacy paradox is to migrate adversarial anonymization frameworks to fully localized environments using local small-scale models (LSMs). However, our empirical investigation reveals that such a naive migration leads to severe utility collapse. As illustrated in Figure 1, the naive migration on Llama3-8B causes destructive over-editing. It indiscriminately strips away non-sensitive context and stylistic nuances, reducing the rich original narrative to a generic and vacuous summary. Besides, it is observed that this collapse persists even when implementing instruction refinement (e.g., outputting "unknown" for inferences with low confidence). This observation demonstrates that heuristic constraints alone cannot rectify the inherent irrationality of the greedy decision-making process.
@@
We argue that this failure cannot be attributed solely to the limited capabilities of LSMs, but instead arises from how greedy adversarial strategies behave under imperfect inferences. From an economic perspective, the anonymization can be viewed as a sequence of decisions balancing Marginal Privacy Gain (MPG) against Marginal Utility Cost (MUC) mas1995microeconomic. By quantifying this ratio as the Marginal Rate of Substitution ($\text{MRS}=\frac{\text{MUC}}{\text{MPG}}$), the primary driver of the utility collapse becomes conceptually clear through this economic lens: Current greedy strategies tend to drift into an economically inefficient state due to LSMs’ hallucinations and the law of diminishing returns, where the model executes destructive edits for non-existent or negligible leaks (see theoretical analysis in Appendix A). In this case, MUC remains positive while MPG approaches zero, pushing MRS to inefficient levels. One concurrent study SEAL kim2025self aims to distill knowledge from a powerful LLM to an LSM through Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) rafailov2023direct to avoid disclosing raw data to third-party APIs. However, our empirical analysis in Section 4.3 shows that it still fails to handle the rationality of anonymization. Besides, SEAL requires additional resources for post-training while high-quality synthetic datasets remain scarce and difficult to construct effectively yukhymenko2024synthetic.
@@
To overcome both the irrationality of naive migration and the data dependency of distillation-based approaches, we propose RLAA (Rational Localized Adversarial Anonymization) [^1] Our code and data are available at https://github.com/SowingG2333/RLAA., a training-free framework designed for fully localized deployment. The core idea of RLAA is to constrain the anonymization process via architectural rationality, thereby compensating for the capability deficits of LSMs without relying on explicit numerical control or parameter fine-tuning. RLAA introduces a novel Attacker–Arbitrator–Anonymizer (A–A–A) architecture, in which the arbitrator acts as a rational gatekeeper between attack feedback and anonymization actions. Moving beyond the paradigm of blindly leaping into modifications, RLAA compels the arbitrator to scrutinize the attacker’s inferences through a validation mechanism inspired by meta-reasoning alazraki2025meta; wen2025reinforcement. This mechanism structurally prevents utility collapse by rejecting destructive edits driven by invalid feedback, and maintains robust privacy protection by focusing the modification budget solely on genuine leaks. Crucially, this design leverages the cognitive asymmetry that verification is less complex and more reliable than generation for LSMs, which may hallucinate during open-ended inference but retain the ability to recognize errors in structured discrimination tasks cobbe2021training; wang2022self; madaan2023self; guan2024language. We conducted a comprehensive evaluation including baseline comparisons, ablation studies and generalization stress tests across several mainstream LLMs (Llama3-8B llama3modelcard, Qwen2.5-7B qwen2.5 and DeepSeek-V3.2-Exp) on the PersonalReddit and reddit-self-disclosure datasets. Additionally, we validated the agent’s rational decision-making process through a quantitative economic efficiency analysis based on the MRS metric. The results empirically demonstrate that RLAA structurally prevents utility collapse and achieves a superior privacy-utility trade-off compared to strong baselines. In summary, the main contributions of this study are as follows:
@@
  • We argue that migrating adversarial anonymization to local environments is crucial for eliminating API-based privacy risks, while a naive migration to LSMs results in utility collapse. The collapse in local adversarial anonymization is identified as a symptom of the economic irrationality inherent in greedy strategies, rather than merely capability deficits.
@@
  • RLAA is proposed as a localized and training-free framework featuring an Attacker-Arbitrator-Anonymizer (A-A-A) architecture. By introducing an arbitrator, it structurally promotes rational decision-making to prevent utility collapse.
@@
  • Extensive experiments demonstrate that RLAA achieves a superior privacy-utility trade-off compared to multiple strong baselines. It achieves Pareto dominance on reddit-self-disclosure and demonstrate its mechanism generalization across different base models.
@@ 2.1 Inference Attacks through LLMs
With increased reasoning capabilities, LLMs have evolved into potent privacy attackers capable of automatically inferring personal attributes from unstructured text wang2025unique. staab24beyond demonstrated that models can accurately deduce fine-grained PII from casual online comments. Extending to high-stakes domains, nyffenegger2024anonymity revealed that LLMs could re-identify individuals in court decisions. Recent audits by panda2025daiq further confirmed that this inference capability is consistent across various model architectures. These studies show that LLMs significantly lower the threshold for privacy violations, enabling re-identification attacks to occur on an unprecedented scale staab24beyond.
@@ 2.2 Text Anonymization
Recent advancements in text anonymization have marked a fundamental evolution from early statistical obfuscation techniques toward sophisticated semantic rewriting mechanisms. Generally, these studies can be divided into two categories:
Traditional Approaches. In early efforts, Differential Privacy (DP) methods abadi2016deep; wu2023privacy; igamberdiev2023dp and representation learning coavoux2018privacy; friedrich2019adversarial served as the standard. However, these methods often compromise textual readability or require training specific encoders, which limits their applicability to open-ended generation tasks. To address utility preservation, shetty2018a4nt introduced A4NT to obfuscate author attributes, while mosallanezhad2019deep utilized deep reinforcement learning to optimize the privacy-utility trade-off. These works laid the foundational mechanism for modern anonymizers.
LLM-driven Adversarial Frameworks. With the rise of generative AI, Feedback-guided Adversarial Anonymization (FgAA) staab2024large established the prevailing paradigm for state-of-the-art anonymization. It systematizes the LLM’s duality, leveraging an attacker model to iteratively critique and refine the anonymizer’s output. Following this paradigm, SEAL kim2025self attempts to distill these capabilities into smaller models through SFT and DPO from multi-round teacher trajectories. However, this mechanism remains constrained by synthetic data scarcity yukhymenko2024synthetic and potential generalization issues. IncogniText frikha2024incognitext employs a misleading strategy, injecting false attributes to randomize attacker inferences rather than simply removing PII. Beyond general-purpose anonymization, RUPTA yang2025robust addresses a complementary problem: optimizing anonymized text for specific downstream tasks (e.g., occupation classification) using specialized task evaluators. While orthogonal to our focus on general semantic preservation, RUPTA shares the common reliance on API-based LLMs. Crucially, a common limitation across these advanced frameworks is their reliance on strong LLMs via API or extensive training pipelines. Moreover, methods following the greedy adversarial paradigm lack rational decision-making mechanisms, making them prone to utility collapse when deployed with smaller local models.
Figure 2: The RLAA Framework. Utilizing an Attacker-Arbitrator-Anonymizer architecture, the arbitrator acts as a rationality gatekeeper. It validates attacker inferences to filter out ghost leaks with negligible privacy benefits, structurally preventing utility collapse caused by irrational greedy strategies.
@@ 3.1 Threat Model
RLAA is designed to defend against two distinct adversaries in the text anonymization task:
Semi-honest Service Provider ($\mathcal{A}_{serv}$). This adversary represents a third-party entity that hosts remote LLM services. It causes the privacy paradox: to utilize superior anonymizers, users must first transmit raw sensitive text $x^{ori}$ to the provider. This transmission exposes users to risks such as unauthorized data retention for model training or commercial profiling smith2023identifying. Distinct from probable inference attacks, we characterize this threat as a deterministic data exposure, where $\mathcal{A}_{serv}$ gains full access to $x^{ori}$ immediately upon submission.
Re-identification Adversary ($\mathcal{A}_{re-id}$). This adversary represents any entity with access to the anonymized output $x^{ano}$. We assume $\mathcal{A}_{re-id}$ employs a powerful LLM $\mathcal{M}_{atk}$ (e.g., DeepSeek-V3.2-Exp in our study) to infer PII. Given an attack prompt $P$, the adversary’s goal is to maximize the accuracy of the inferred attribute-value tuples $\{(a_{j},v_{j})\}$, where $a_{j}$ denotes a specific private attribute category (e.g., location) and $v_{j}$ represents its corresponding inferred value (e.g., "Paris, France"):
$\{(a_{j},v_{j})\}\leftarrow\mathcal{M}_{atk}(P(x^{ano}))$ (1)
@@ 3.2 Problem Formulation
The text anonymization task can be formalized as a constrained transformation problem. Let $\mathcal{X}$ be the space of all text sequences. Given an original text $x^{(0)}\in\mathcal{X}$, let $\mathcal{A}_{priv}=\{(a_{k},v_{k}^{*})\}_{k=1}^{K}$ be the set of "private attribute-true value" pairs contained in $x^{(0)}$, where $a_{k}$ represents the attribute category (e.g., age) and $v_{k}^{*}$ denotes its true value.
@@
$P_{atk}(x)=\frac{1}{K}\sum_{k=1}^{K}\mathbb{I}\left(\hat{v}_{k}=v_{k}^{*}\mid(a_{k},\hat{v}_{k})\in\mathcal{M}_{atk}(P(x))\right)$ (2)
where $\mathbb{I}(\cdot)$ is the indicator function checking if the inferred value $\hat{v}_{k}$ matches the ground truth $v_{k}^{*}$.
Simultaneously, let $U(x^{(0)},x):\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be a utility function measuring semantic preservation, where a higher value indicates better utility. The ideal objective is to find an optimum $x^{*}$ such that private attributes are protected while utility is maximized:
$\max\quad U(x^{(0)},x)\quad\text{s.t.}\quad P_{atk}(x)\leq\delta$ (3)
where $\delta$ represents the upper bound of acceptable privacy risk.
Migrating the framework to a local environment naturally eliminates the service provider threat $\mathcal{A}_{serv}$. However, achieving the objective in Eq. 3 is non-trivial. Existing adversarial frameworks typically approximate the optimum $x^{*}$ via an iterative sequence $x^{(0)}\to x^{(1)}\dots\to x^{(T)}$, where each step attempts to reduce privacy risk. Our empirical investigation reveals a critical challenge: directly applying greedy adversarial iterations with LSMs in this process often leads to utility collapse. We hypothesize that this failure arises because naive greedy agents lack a mechanism to evaluate the cost-effectiveness of privacy gains. To formally address this, we reframe this iterative process over steps $t=1,2\dots T$ through an economic lens, viewing each anonymization operation as a transaction. The following marginal metrics are defined to quantify the rationality of each update.
@@
$\Delta P_{t}=P_{atk}(x^{(t-1)})-P_{atk}(x^{(t)})$ (4)
where $x^{(t-1)}$ represents the intermediate text generated at the previous iteration.
@@
The instantaneous price of privacy preservation, representing the utility cost per unit of privacy gained:
@@
From this perspective, a rational agent should ideally align with the principle that $\text{MRS}_{t}\leq\lambda$, where $\lambda$ represents the maximum rational utility cost per unit of privacy. Naive greedy agents inherently tends to violate this condition and drift into irrational states of deadweight loss where $\text{MRS}_{t}\to\infty$. Guided by this insight, RLAA is proposed to implicitly enforce this budget constraint $\lambda$ through architectural rationality, thereby structurally preventing utility collapse. Detailed theoretical analysis is shown in Appendix A.